OpenStudy (anonymous):

Which typing time is the fastest? 320 words in 8 minutes 600 words in 12 minutes 350 words in 10 minutes 225 words in 5 minutes

OpenStudy (anonymous):

it's b

OpenStudy (anonymous):

600 words in 12 min

OpenStudy (anonymous):

IT WOULD HAVE TO BE 600 WORDS IN 12MINUTES

OpenStudy (charlotte123):

I accept the challange

OpenStudy (charlotte123):

oh and to find out the answer (or check it), u just take the number of words and divide it by the number of minutes :)

OpenStudy (anonymous):

UR CORRECT

OpenStudy (charlotte123):

of course I am HUHEHEUEHUE

OpenStudy (charlotte123):

wink wink double pink

OpenStudy (texaschic101):

divide each one of them.. 320/8 = 600/12 = and so on....the highest number is the fastest

OpenStudy (221emily):

you have to getting this from a book your the funniest person I have ever met!!! lol

OpenStudy (charlotte123):

@221emily I get it from a book called, Charlottes Book of Puns

OpenStudy (charlotte123):

was born with it hun, its all in the face

OpenStudy (charlotte123):

HUEHEUHEUEHUE

OpenStudy (221emily):

lol I got it

OpenStudy (anonymous):

hahahahhaha right one

OpenStudy (anonymous):

o-o meep thanks everyone!

OpenStudy (charlotte123):

@221emily I'll give ya one for free *hands u copy* take care Em :D make people laugh until they shoot strawberry's from their eyes!

OpenStudy (charlotte123):

@foreverlost was my pleasure, HUEHUHEUE

OpenStudy (anonymous):

ya anytime

OpenStudy (221emily):

ok yeah defiantly do that